OpenStudy (anonymous):

I have a few questions I need help with involving limits and continuous functions. Since they all have equations it will take me a minute or two to type all the equations up and post!

OpenStudy (anonymous):

find the value of c such that \[\lim_{x \rightarrow c ^{+}} \frac{ 6 }{ (x-2)(7-x)}=-\infty\]

OpenStudy (perl):

why dont you graph f(x) = 6 / [ (x-2)(7-x)] first

OpenStudy (perl):

c = 7 works

OpenStudy (aum):

The expression becomes infinity when the denominator gets to zero. There are two values of x that will make the denominator zero. Which value when approached from the left will yield NEGATIVE infinity?

OpenStudy (aum):

I meant "approached from the right"

OpenStudy (perl):